# Connection between memory performance and optical absorption in quantum reservoir computing

### v1

@@ Acknowledgements
Appendix A Absorption spectra
To gain a better insight regarding the structure of the absorption curves in Fig. 4 in the main text, we here show three exemplary absorption spectra for the input strength $s=1$ and three different qubit decays in Fig. 5. For a small decay ($\gamma=10^{-2}$), we find sharp peaks centered around the zero-frequency in the absorption spectrum, which are a result of the external pumping and change in position for different input signals. Upon increasing the dissipation strength, the typical line broadening occurs and allows for the absorption of information in the pump frequency of $\omega=0$ – it is around this dissipation that the absorption, and with it the STMC, is maximized. Elevating the dissipation even further to $\gamma=100$ leads to an almost completely flat spectrum, as the line broadening washes away all structure in the spectrum.
Appendix B Infinite shot STMC
As discussed in the main text, Fig. 3 depicts the STMC for finite shot noise corresponding to $10^{6}$ measurements. In the infinite shot case as shown in Fig. 6, however, we observe a slightly different behavior of the STMC for larger decays; even the most minute deviations from the zero-input steady state can be utilized by the noiseless reservoir computer to extract the current input, resulting in a linear STMC of almost exactly one. We find that higher orders of the STMC do not benefit to the same degree from the absence of noise and still drop close to zero for large decays.

### v2

@@ Acknowledgements
Appendix A SUPPLEMENTARY MATERIAL
Appendix B Absorption spectra
To gain a better insight regarding the structure of the absorption curves in Fig. 4 in the main text, we here show three exemplary absorption spectra for the input strength $s=1$ and three different qubit decays in Fig. 5. For a small decay ($\gamma=10^{-2}$), we find sharp peaks centered around the zero-frequency in the absorption spectrum, which are a result of the external pumping and change in position for different input signals. Upon increasing the dissipation strength, the typical line broadening occurs and allows for the absorption of information encoded in the pump frequency of $\omega=0$ – it is around this dissipation that the absorption, and with it the STMC, is maximized. Elevating the dissipation even further to $\gamma=100$ leads to an almost completely flat spectrum, as the line broadening washes away all structure in the spectrum.
Furthermore, the absorption curves in Fig. 4 exhibit shoulders around the maximum absorption. To gain insight into this observation, in Fig. 6, we take as an example a single 3-qubit reservoir with a coupling strength of $J_{0}=0.5$ exposed to a signal of amplitude $s=1$ for a range of qubit decays. For the smallest decay (here, $\gamma=1$), the resonant absorption (i.e. absorption at $\omega=0$ in the rotating frame) of the TFIM lies in between two peaks. As the peaks merge asymmetrically for increasing decays, the resonant absorption behaves non-monotonically until the peaks are fully merged at around $\gamma=5$, after which the absorption decreases steadily. This behavior gives rise to the shoulders appearing in Fig. 4 and strongly depends on the specific coupling and signal strength. It cannot trivially be bypassed by changing the drive frequency, as this also changes the peak positions.
Figure 6: Absorption spectra for a 3-qubit reservoir computer with a coupling strength $J_{0}=0.5$ exposed to a signal of amplitude $s=1$ for a range of qubit decays. The $\alpha$-axis is cropped to make the changes of the resonant absorption (blue dots) more visible. We observe a non-monotonous behavior of the absorption with respect to the decay rate.
Appendix C Infinite shot STMC
As discussed in the main text, Fig. 3 depicts the STMC for finite shot noise corresponding to $10^{6}$ measurements. In the infinite-shot case, which is shown here in Fig. 7, we observe a slightly different behavior of the STMC for larger decays: Even the most minute deviations from the zero-input steady state can be utilized by an ideal, completely noise-free reservoir computer to extract the current input, resulting in a linear STMC of almost exactly one. We find that higher orders of the STMC do not benefit to the same degree from the absence of noise and still drop close to zero for large decays.
@@ Appendix C Infinite shot STMC
Appendix D STMC to absorption relation generality
In Fig. 8, we show the same simulations as carried out in the main text for 3 qubits in Figs. 3 and 4, but for different coupling strengths $J_{0}\in\{0.25,1.0\}$ and two 4-qubit reservoirs with all-to-all and ring topology, respectively. The results remain qualitatively the same as for the systems considered in the main text, indicating the generality of the revealed connection between absorption and memory capacity also for larger systems, different topologies, and coupling strengths.
Figure 8: The STMC and absorption for different coupling strengths, system sizes, and topologies. We observe the same behavior as for the systems considered in the main text.
QRC
quantum reservoir computing
QELM
quantum extreme learning machine
RC
reservoir computing
TFIM
transverse-field Ising model
NISQ
noisy intermediate-scale quantum
GKSL
Gorini–Kossakowski–Sudarshan–Lindblad
IPC
information processing capacity
STMC
short-term memory capacity
RNN
recurrent neural network